Find the slope of the line containing each given pair of points. If the slope is undefined, state this.
(-1,4) and (5,-8)

I have know idea how to do this problem, could someone please help me.
Thank you in advance

Answer by stanbon(75887)

Find the slope of the line containing each given pair of points. If the slope is undefined, state this.
(-1,4) and (5,-8)
-----------------
slope = (change in y)/(change in x)
----
Your Problem:
m = (-8-4)/(5--1) = -12/6 = -2
=================================
Cheers,
Stan H.